# RoMedQA: The First Benchmark for Romanian Medical Question Answering

### v1

@@ Abstract
Question answering (QA) is an actively studied topic, being a core natural language processing (NLP) task that needs to be addressed before achieving Artificial General Intelligence (AGI). However, the lack of QA datasets in specific domains and languages hinders the development of robust AI models able to generalize across various domains and languages. To this end, we introduce RoMedQA, the first Romanian QA benchmark for the medical domain, alongside a comprehensive evaluation of state-of-the-art large language models (LLMs). We construct a high-quality and large-scale dataset comprising 102,646 QA pairs related to cancer patients. The questions regard medical case summaries of 1,011 patients, requiring either keyword extraction or reasoning to be answered correctly. RoMedQA is the result of a time-consuming manual annotation process carried out by seven physicians specialized in oncology or radiotherapy, who spent a total of about 2,100 work hours to generate the QA pairs. We experiment with four LLMs from distinct families of models on RoMedQA. Each model is employed in two scenarios, namely one based on zero-shot prompting and one based on supervised fine-tuning. Our results show that fine-tuned models significantly outperform their zero-shot counterparts, clearly indicating that pretrained models fail to generalize on RoMedQA. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian. We publicly release our dataset and code at https://github.com/ana-rogoz/RoMedQA.
@@ 1 Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) Pal and Sankarasubbu (2024); Touvron et al. (2023); Jiang et al. (2023); Taori et al. (2023); Köpf et al. (2023), which are able to perform remarkably well across various QA tasks Brown et al. (2020); Chowdhery et al. (2023); Rajpurkar et al. (2016); Hendrycks et al. (2021); Achiam et al. (2023); Singhal et al. (2023, 2025), QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents. While large-scale datasets such as PubMedQA Jin et al. (2019), emrQA Pampari et al. (2018), and MedQA Jin et al. (2021) have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. To the best of our knowledge, for the Romanian language, there are no prior efforts in the realm of QA resources for the medical domain.
To address this gap, we introduce RoMedQA, the first benchmark for Romanian Medical Question Answering. As illustrated in Figure 1, the benchmark creation involves two phases, (i) manual data collection and annotation, and (ii) model benchmarking. The dataset comprises 102,646 high-quality QA pairs from real-world clinical records of 1,011 oncology patients (796 patients with breast cancer and 215 patients with lung cancer). The QA pairs are the results of a manual annotation process carried out by physicians specialized in oncology and radiotherapy, who collectively spent nearly 2,100 hours to generate the pairs, ensuring both linguistic quality and medical correctness. RoMedQA includes 76,416 QA pairs about breast cancer patients and 26,230 about lung cancer patients, with questions grounded in medical case summaries (epicrises).
@@
RoMedQA (ours) Romanian Medical 102,646
Table 1: Comparison between RoMedQA and existing QA datasets across three dimensions, namely language coverage, domain focus, and dataset size. Existing QA datasets range from general-purpose English resources to multilingual datasets and domain-specific collections in areas such as legal, technical and medical fields. RoMedQA is the first benchmark for medical QA in Romanian.
To prevent data leakage across splits and support fair model development and evaluation, we provide official training, validation, and test splits at the patient level, i.e. a medical record can appear in only one of the splits. We benchmark four models on RoMedQA, namely two LLMs specialized on Romanian (RoLLaMA2-7B and RoMistral-7B) Masala et al. (2024), a long-context LLM (Phi-4-mini-instruct) Abouelenin et al. (2025), and an LLM tuned on biomedical data (LLaMA3-OpenBioLLM-8B) Pal and Sankarasubbu (2024). Each model is assessed in both zero-shot and supervised fine-tuning setups, while testing various prompt formats. Our results show that all LLMs significantly benefit from fine-tuning, whereas zero-shot variants perform poorly. These findings emphasize the critical importance of both domain and language adaptation for reliable clinical QA in Romanian. Among all evaluated models, the fine-tuned Phi-4-mini-instruct achieves the best performance, significantly outperforming other approaches. However, the best-performing model achieves an F1 score of only $0.667$, suggesting that RoMedQA is a challenging dataset, highlighting the need for the development of more robust models, capable of better generalizing to specialized domains and low-resource languages.
@@
  • We introduce RoMedQA, the first and only large-scale medical QA dataset in Romanian, comprising 102,646 QA instances obtained with the help of trained physicians.
@@ 2 Related Work
In contrast to the significant advances in English medical QA, the landscape of non-English medical QA datasets remains scarce. To the best of our knowledge, we are the first to introduce a Romanian QA dataset specifically tailored for the medical domain. RoMedQA contains original clinically-grounded content rather than translations, addressing a critical gap in current literature.
Romanian QA datasets. For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset Crăciun et al. (2025), which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset Nicolae and Tufiş (2021), which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark Dumitrescu et al. (2021), which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD Artetxe et al. (2020) for Romanian, consisting of translated versions of English questions. None of the existing datasets for Romanian QA is tailored for the medical domain. Moreover, existing datasets are typically small (containing between 1,000 and 11,000 QA pairs). With 102,646 QA pairs, RoMedQA is one order of magnitude larger than the largest dataset for Romanian QA.
RoMedQA vs. QA dataset landscape. In Table 1, we present a comprehensive comparison between RoMedQA and existing QA datasets across different languages and domains. RoMedQA fills in a gap in the landscape of medical QA resources by providing the first large-scale Romanian medical QA dataset with 102,646 question-answer pairs focused on oncology. While other Romanian QA datasets exist for other domains, none of them specifically addresses the medical domain with clinically-relevant content in Romanian. This positions RoMedQA as the first and only resource for advancing medical NLP capabilities for Romanian, an underrepresented language in the QA field.
@@
Table 2: A summary of the number of patients and corresponding QA pairs included in the RoMedQA benchmark across two types of cancer, breast and lung. The splits are performed at the patient level, into training (70%), validation (15%), and test (15%), to prevent information leakage across splits. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across both cancer types for each split. RoMedQA comprises a total 102,646 QA pairs for 1,011 patients.
@@ 3.1 Data Collection
RoMedQA is obtained through the manual extraction of information from medical documents, namely medical summaries of patients registered in the Oncology Department of Colţea Clinical Hospital in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing RoMedQA. In this group, two experienced physicians, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnostics, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer.
@@ 3.3 Data Partitioning and Statistics
We present the distribution of our dataset across training, validation, and test sets in Table 2. For the breast cancer subset, our split results in 557 patients allocated for training, 119 for validation, and 120 for testing. In terms of QA pairs, the split distributes 53,472 samples for training, 11,424 for validation, and 11,520 for testing, respectively. Similarly, the lung cancer subset is divided into 150 training, 32 validation, and 33 test patients, yielding 18,300, 3,904, and 4,026 QA pairs, respectively. The splits are independently generated for each cancer type. We next combine the corresponding partitions to create our final unified dataset, called RoMedQA. In total, RoMedQA comprises 1,011 unique patients (796 suffering from breast cancer, and 215 from lung cancer) and 102,646 question-answer pairs, providing a substantial resource for training and evaluating Romanian medical QA systems.
In Table 3, we report statistics about the number of tokens in epicrises, questions and answers included RoMedQA. Epicrises range from 207 to 34,247 tokens, with a mean value of 7,171 tokens, where shorter documents typically correspond to patients under investigation or at early treatment stages, while longer ones capture extensive treatment histories with multiple interventions. This variation can pose challenges for some models with limited context windows. While epicrises are typically lengthy, questions and answers are much shorter. Questions maintain relative consistency, between 7 and 30 tokens, with a mean value of 15 tokens, reflecting the consistency of clinical query formulations. The answers are represented by predominantly brief texts with a mean of 4 tokens, varying in length from one to 442 tokens. Longer responses are typical for questions involving complex medical reasoning.
@@
Table 3: A summary of the minimum, maximum, average and total number of tokens for the three kinds of data (epicrisis, question and answer) available in RoMedQA.
Figure 2: Distribution of question types in the RoMedQA dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 102,646 QA pairs. Best viewed in color.
We categorize the questions into three types: (i) binary (yes/no); (ii) extractive, where the answer appears explicitly in the clinical context (epicrisis); and (iii) reasoning, where the answer must be inferred from multiple clues within the epicrisis. The distribution of these categories in RoMedQA is shown in Figure 2. Within the binary subset, the label distribution is moderately skewed, with roughly two thirds of answers being “Nu” (“No”) and one third being “Da” (“Yes”). The slightly skewed distribution resulted naturally from the data collection process.
@@ 4 Methods
All four LLMs are employed in two setups, zero-shot prompting and end-to-end fine-tuning. Due to the large number of parameters in the chosen models, we adopt a parameter-efficient fine-tuning strategy based on Low-Rank Adaptation (LoRA) Hu et al. (2022) specialized on the causal language modeling task. By comparing the results obtained in the two setups, we can precisely contextualize the impact of fine-tuning on RoMedQA. In addition, we also test two alternative prompts, where the order between context (epicrisis) and question is interchanged. Finally, to determine the impact of zero-shot prompting, we include a random token selector and a majority answer model as baselines for the QA task. Next, we describe all evaluated methods in detail.
@@ 4.1 Romanian-Adapted LLMs
We start from two instruction-tuned language models specifically adapted for Romanian, RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Both models are part of the recent efforts to build high-quality Romanian LLMs by further pretraining on Romanian corpora and applying instruction tuning with Romanian datasets, such as RoAlpaca Taori et al. (2023); Masala et al. (2024), RoOpenAssistant Köpf et al. (2023); Masala et al. (2024) and RoOrca Mukherjee et al. (2023); Masala et al. (2024). The two LLMs have been optimized for Romanian morphology and syntax using tokenizers adapted to the language. They have shown strong performance on Romanian natural language understanding and generation tasks Masala et al. (2024). In our study, we further fine-tune both models on RoMedQA and compare their performance under different prompt structures and lengths.
@@ 5 Experiments
We conduct experiments with the chosen methods on the RoMedQA dataset, using our official split. We start by tuning various hyperparameters involved in the fine-tuning process, aiming to find the optimal hyperparameter setup for all LLMs. We next evaluate two different prompt formats, to determine which prompt format leads to higher performance. For the best performing prompt structure, we conduct several experiments across multiple LLMs, using various prompt lengths. The objectives of our experiments are: (i) to capture how Romanian-adapted models compare to general-purpose and domain-specific alternatives, (ii) to determine how fine-tuning performs against zero-shot inference, (iii) to establish how fine-tuned or zero-shot LLMs compare against trivial baselines, and (iv) to investigate how the context length influences model performance, particularly for models with extended context capabilities.
@@ 5.1 Evaluation
BLEU. Because RoMedQA contains answers with more than one token, we consider the BLEU score Papineni et al. (2002), which measures n-gram overlap to assess fluency and lexical similarity. When the predicted answer includes only a subset of the tokens from the ground-truth answer, BLEU can provide a representative measure by considering partial n-gram matches.
@@
Table 5: Results of multiple LLMs based on various configurations vs. two trivial baselines on RoMedQA. For each LLM, there is a zero-shot version and several fine-tuned versions, with various prompt lengths. F1, Exact Match (EM), BLEU, and METEOR scores are reported on both validation and test sets. Fine-tuned models consistently outperform zero-shot versions. The best results for each LLM are highlighted in bold green.

### v2

# MedQARo: A Large-Scale Benchmark for Medical Question Answering in Romanian

@@ Abstract
Question answering (QA) is an actively studied topic, being a core natural language processing (NLP) task that needs to be addressed before achieving Artificial General Intelligence (AGI). However, the lack of QA datasets in specific domains and languages hinders the development of robust AI models able to generalize across various domains and languages. To this end, we introduce MedQARo, the first large-scale medical QA benchmark in Romanian, alongside a comprehensive evaluation of state-of-the-art large language models (LLMs). We construct a high-quality and large-scale dataset comprising 102,646 QA pairs related to cancer patients. The questions regard medical case summaries of 1,011 patients, requiring either keyword extraction or reasoning to be answered correctly. MedQARo is the result of a time-consuming manual annotation process carried out by seven physicians specialized in oncology or radiotherapy, who spent a total of about 2,100 work hours to generate the QA pairs. We experiment with four LLMs from distinct families of models on MedQARo. Each model is employed in two scenarios, namely one based on zero-shot prompting and one based on supervised fine-tuning. Our results show that fine-tuned models significantly outperform their zero-shot counterparts, clearly indicating that pretrained models fail to generalize on MedQARo. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian. We publicly release our dataset and code at https://github.com/ana-rogoz/MedQARo.
@@ 1 Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) Pal and Sankarasubbu (2024); Touvron et al. (2023); Jiang et al. (2023); Taori et al. (2023); Köpf et al. (2023), which are able to perform remarkably well across various QA tasks Brown et al. (2020); Chowdhery et al. (2023); Rajpurkar et al. (2016); Hendrycks et al. (2021); Achiam et al. (2023); Singhal et al. (2023, 2025), QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents. While large-scale datasets such as PubMedQA Jin et al. (2019), emrQA Pampari et al. (2018), and MedQA Jin et al. (2021) have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. To the best of our knowledge, for the Romanian language, there are no prior efforts in the realm of large-scale QA resources for the medical domain.
To address this gap, we introduce MedQARo, the first large-scale benchmark for Medical Question Answering in Romanian (MedQARo). As illustrated in Figure 1, the benchmark creation involves two phases, (i) manual data collection and annotation, and (ii) model benchmarking. The dataset comprises 102,646 high-quality QA pairs from real-world clinical records of 1,011 oncology patients (796 patients with breast cancer and 215 patients with lung cancer). The QA pairs are the results of a manual annotation process carried out by physicians specialized in oncology and radiotherapy, who collectively spent nearly 2,100 hours to generate the pairs, ensuring both linguistic quality and medical correctness. MedQARo includes 76,416 QA pairs about breast cancer patients and 26,230 about lung cancer patients, with questions grounded in medical case summaries (epicrises).
@@
RoMedQA Dima et al. (2024) Romanian Biology exams 4,127
MedQARo (ours) Romanian Medical cases 102,646
Table 1: Comparison between MedQARo and existing QA datasets across three dimensions, namely language coverage, domain focus, and dataset size. Existing QA datasets range from general-purpose English resources to multilingual datasets and domain-specific collections in areas such as legal, technical and medical fields. MedQARo is the first large-scale benchmark for medical QA in Romanian.
To prevent data leakage across splits and support fair model development and evaluation, we provide official training, validation, and test splits at the patient level, i.e. a medical record can appear in only one of the splits. We benchmark four models on MedQARo, namely two LLMs specialized on Romanian (RoLLaMA2-7B and RoMistral-7B) Masala et al. (2024), a long-context LLM (Phi-4-mini-instruct) Abouelenin et al. (2025), and an LLM tuned on biomedical data (LLaMA3-OpenBioLLM-8B) Pal and Sankarasubbu (2024). Each model is assessed in both zero-shot and supervised fine-tuning setups, while testing various prompt formats. Our results show that all LLMs significantly benefit from fine-tuning, whereas zero-shot variants perform poorly. These findings emphasize the critical importance of both domain and language adaptation for reliable clinical QA in Romanian. Among all evaluated models, the fine-tuned Phi-4-mini-instruct achieves the best performance, significantly outperforming other approaches. However, the best-performing model achieves an F1 score of only $0.667$, suggesting that MedQARo is a challenging dataset, highlighting the need for the development of more robust models, capable of better generalizing to specialized domains and low-resource languages.
@@
  • We introduce MedQARo, the first large-scale medical QA dataset in Romanian, comprising 102,646 QA instances obtained with the help of trained physicians.
@@ 2 Related Work
In contrast to the significant advances in English medical QA, the landscape of non-English medical QA datasets remains scarce. To the best of our knowledge, we are the first to introduce a large-scale QA dataset specifically tailored for the medical domain in Romanian. MedQARo contains original clinically-grounded content rather than translations, addressing a critical gap in current literature.
Romanian QA datasets. For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset Crăciun et al. (2025), which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset Nicolae and Tufiş (2021), which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark Dumitrescu et al. (2021), which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD Artetxe et al. (2020) for Romanian, consisting of translated versions of English questions. For advanced biology, Dima et al. (2024) published a small-scale dataset comprising 4,127 single-choice questions. Existing datasets for Romanian QA are typically small (containing between 1,000 and 11,000 QA pairs). With 102,646 QA pairs, MedQARo is one order of magnitude larger than the largest dataset for Romanian QA.
MedQARo vs. QA dataset landscape. In Table 1, we present a comprehensive comparison between MedQARo and existing QA datasets across different languages and domains. MedQARo fills in a gap in the landscape of medical QA resources by providing the first large-scale Romanian medical QA dataset with 102,646 question-answer pairs focused on oncology. While other Romanian QA datasets exist, none of them includes extractive and reasoning questions based on medical case summaries. This positions MedQARo as the first large-scale resource for advancing medical NLP capabilities for Romanian, an underrepresented language in the QA field.
@@
Table 2: A summary of the number of patients and corresponding QA pairs included in the MedQARo benchmark across two types of cancer, breast and lung. The splits are performed at the patient level, into training (70%), validation (15%), and test (15%), to prevent information leakage across splits. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across both cancer types for each split. MedQARo comprises a total 102,646 QA pairs for 1,011 patients.
@@ 3.1 Data Collection
MedQARo is obtained through the manual extraction of information from medical documents, namely medical summaries of patients registered in the Oncology Department of Colţea Clinical Hospital in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing MedQARo. In this group, two experienced physicians, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnostics, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer.
@@ 3.3 Data Partitioning and Statistics
We present the distribution of our dataset across training, validation, and test sets in Table 2. For the breast cancer subset, our split results in 557 patients allocated for training, 119 for validation, and 120 for testing. In terms of QA pairs, the split distributes 53,472 samples for training, 11,424 for validation, and 11,520 for testing, respectively. Similarly, the lung cancer subset is divided into 150 training, 32 validation, and 33 test patients, yielding 18,300, 3,904, and 4,026 QA pairs, respectively. The splits are independently generated for each cancer type. We next combine the corresponding partitions to create our final unified dataset, called MedQARo. In total, MedQARo comprises 1,011 unique patients (796 suffering from breast cancer, and 215 from lung cancer) and 102,646 question-answer pairs, providing a substantial resource for training and evaluating Romanian medical QA systems.
In Table 3, we report statistics about the number of tokens in epicrises, questions and answers included MedQARo. Epicrises range from 207 to 34,247 tokens, with a mean value of 7,171 tokens, where shorter documents typically correspond to patients under investigation or at early treatment stages, while longer ones capture extensive treatment histories with multiple interventions. This variation can pose challenges for some models with limited context windows. While epicrises are typically lengthy, questions and answers are much shorter. Questions maintain relative consistency, between 7 and 30 tokens, with a mean value of 15 tokens, reflecting the consistency of clinical query formulations. The answers are represented by predominantly brief texts with a mean of 4 tokens, varying in length from one to 442 tokens. Longer responses are typical for questions involving complex medical reasoning.
@@
Table 3: A summary of the minimum, maximum, average and total number of tokens for the three kinds of data (epicrisis, question and answer) available in MedQARo.
We categorize the questions into three types: (i) binary (yes/no); (ii) extractive, where the answer appears explicitly in the clinical context (epicrisis); and (iii) reasoning, where the answer must be inferred from multiple clues within the epicrisis. The distribution of these categories in MedQARo is shown in Figure 2. Within the binary subset, the label distribution is moderately skewed, with roughly two thirds of answers being “Nu” (“No”) and one third being “Da” (“Yes”). The slightly skewed distribution resulted naturally from the data collection process.
@@ 4 Methods
All four LLMs are employed in two setups, zero-shot prompting and end-to-end fine-tuning. Due to the large number of parameters in the chosen models, we adopt a parameter-efficient fine-tuning strategy based on Low-Rank Adaptation (LoRA) Hu et al. (2022) specialized on the causal language modeling task. By comparing the results obtained in the two setups, we can precisely contextualize the impact of fine-tuning on MedQARo. In addition, we also test two alternative prompts, where the order between context (epicrisis) and question is interchanged. Finally, to determine the impact of zero-shot prompting, we include a random token selector and a majority answer model as baselines for the QA task. Next, we describe all evaluated methods in detail.
Figure 2: Distribution of question types in the MedQARo dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 102,646 QA pairs. Best viewed in color.
@@ 4.1 Romanian-Adapted LLMs
We start from two instruction-tuned language models specifically adapted for Romanian, RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Both models are part of the recent efforts to build high-quality Romanian LLMs by further pretraining on Romanian corpora and applying instruction tuning with Romanian datasets, such as RoAlpaca Taori et al. (2023); Masala et al. (2024), RoOpenAssistant Köpf et al. (2023); Masala et al. (2024) and RoOrca Mukherjee et al. (2023); Masala et al. (2024). The two LLMs have been optimized for Romanian morphology and syntax using tokenizers adapted to the language. They have shown strong performance on Romanian natural language understanding and generation tasks Masala et al. (2024). In our study, we further fine-tune both models on MedQARo and compare their performance under different prompt structures and lengths.
@@ 5 Experiments
We conduct experiments with the chosen methods on the MedQARo dataset, using our official split. We start by tuning various hyperparameters involved in the fine-tuning process, aiming to find the optimal hyperparameter setup for all LLMs. We next evaluate two different prompt formats, to determine which prompt format leads to higher performance. For the best performing prompt structure, we conduct several experiments across multiple LLMs, using various prompt lengths. The objectives of our experiments are: (i) to capture how Romanian-adapted models compare to general-purpose and domain-specific alternatives, (ii) to determine how fine-tuning performs against zero-shot inference, (iii) to establish how fine-tuned or zero-shot LLMs compare against trivial baselines, and (iv) to investigate how the context length influences model performance, particularly for models with extended context capabilities.
@@ 5.1 Evaluation
BLEU. Because MedQARo contains answers with more than one token, we consider the BLEU score Papineni et al. (2002), which measures n-gram overlap to assess fluency and lexical similarity. When the predicted answer includes only a subset of the tokens from the ground-truth answer, BLEU can provide a representative measure by considering partial n-gram matches.
@@
Table 5: Results of multiple LLMs based on various configurations vs. two trivial baselines on MedQARo. For each LLM, there is a zero-shot version and several fine-tuned versions, with various prompt lengths. F1, Exact Match (EM), BLEU, and METEOR scores are reported on both validation and test sets. Fine-tuned models consistently outperform zero-shot versions. The best results for each LLM are highlighted in bold green.